In a water refilling station, one pipe can fill a tank in 30 min. while another pipe can fill it in 40 min. A third pipe can empty the same tank in 20 min. If all pipes are opened at the same time. in how many minutes will the tan be filled?
Answer by Alan3354(69443)

In a water refilling station, one pipe can fill a tank in 30 min. while another pipe can fill it in 40 min. A third pipe can empty the same tank in 20 min. If all pipes are opened at the same time. in how many minutes will the tan be filled?
-----------
one pipe can fill a tank in 30 min.
Adds 1/30 of a tank per minute.
----------------
while another pipe can fill it in 40 min.
Adds 1/40 of a tank per minute.
----------------
A third pipe can empty the same tank in 20 min
Subtracts 1/20 of a tank per minute.
====================
Each minute, 1/30 + 1/40 - 1/20 of the tank
= 4/120 + 3/120 - 6/120 = 1/120
--> 120 minutes